$SPAD/src/input ovar.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool ovar.output
)set message test on
)set message auto off
)clear all
--S 1 of 5
ls:List Symbol:=['x,'a,'z]
--R
--R
--R   (1)  [x,a,z]
--R                                                           Type: List(Symbol)
--E 1
--S 2 of 5
Z:=OVAR ls
--R
--R
--R   (2)  OrderedVariableList([x,a,z])
--R                                                                 Type: Domain
--E 2
--S 3 of 5
size()$Z
--R
--R
--R   (3)  3
--R                                                     Type: NonNegativeInteger
--E 3
--S 4 of 5
lv:=[index(i::PI)$Z for i in 1..size()$Z]
--R
--I   Compiling function G1409 with type Integer -> Boolean
--I   Compiling function G1573 with type NonNegativeInteger -> Boolean
--R
--R   (4)  [x,a,z]
--R                                     Type: List(OrderedVariableList([x,a,z]))
--E 4
--S 5 of 5

sorted?(>,lv)
--R
--R
--R   (5)  true
--R                                                                Type: Boolean
--E 5
)spool
)lisp (bye)

———-

References^{[1] nothing}

5